Darth Maul is doing some mechanical work on a new engine for his landspeeder. Engine oil (assume he is using SAE 10 ... see table below) passes through a 1.30 mm diameter tube in his newly created engine. The tube is 5.4 cm long. What pressure difference is needed to maintain a flow rate of 5.2 mL/min? (Note: 1 mL = 1 cm3)
Pa

The coefficients of viscosity for various fluid

Coefficients of Viscosity Fluid Coefficient (tempe
Answer #1

we know

Poiseuille's Equation

where \Delta x=length

rest all symbol we know

so we get

Q=5.2mL/min=8.667*10^(-8)

r=0.65mm =0.65*10^(-3) m

viscosity=200*10^(-3)

so we got p=13352.5 pa